# Neural Concept Binder

### v1

@@ Abstract
The challenge in object-based visual reasoning lies in generating descriptive yet distinct concept representations. Moreover, doing this in an unsupervised fashion requires human users to understand a model’s learned concepts and potentially revise false concepts. In addressing this challenge, we introduce the Neural Concept Binder, a new framework for deriving discrete concept representations resulting in what we term “concept-slot encodings”. These encodings leverage both "soft binding" via object-centric block-slot encodings and "hard binding" via retrieval-based inference. The Neural Concept Binder facilitates straightforward concept inspection and direct integration of external knowledge, such as human input or insights from other AI models like GPT-4. Additionally, we demonstrate that incorporating the hard binding mechanism does not compromise performance; instead, it enables seamless integration into both neural and symbolic modules for intricate reasoning tasks, as evidenced by evaluations on our newly introduced CLEVR-Sudoku dataset. [^1] Code and data available here.{NoHyper}These authors share equal contribution.{NoHyper}Correspondence to: Wolfgang Stammer ¡wolfgang.stammer@cs.tu-darmstadt.de¿.
@@ 1 Introduction
Figure 1: Unsupervised learning of concepts for visual reasoning. Models that learn concepts from unlabeled data require inspectable and revisable concept representations.
An essential aspect of visual reasoning is obtaining a proper conceptual understanding of the world by learning visual concepts and processing these into a suitable representation (cf. Fig.  1). The majority of current machine learning (ML) approaches that focus on visual concept-based processing utilize forms of supervised [28, 59, 26], weakly-supervised [34, 60, 41, 54, 4, 72] or text-guided [23] learning of concepts. These approaches all require some form of additional (prior) knowledge about the relevant domain. An attractive alternative, though much more challenging, is to learn concepts in an unsupervised fashion. This comes with several challenges: (i) learning an expressive concept representation without concept supervision is intrinsically difficult [33], and (ii) there is no guarantee that learned concepts align with general domain knowledge [30, 75, 5] and can therefore be utilized for complex downstream tasks. Moreover, (iii) to trust that the learned concept representations are reliable for high stakes scenarios [10], it is necessary to make the model’s concept representations human-inspectable and -revisable [59, 60, 25] (cf. Fig.  1).
@@
This work proposes the Neural Concept Binder (NCB) framework to learn expressive yet inspectable and revisable concepts from unlabeled data. NCB combines continuous encodings, obtained via block-slot-based soft-binding, with discrete concept representations, obtained via retrieval-based hard-binding. Thus, to tackle the challenges of unsupervised concept learning, NCB harnesses the advantages of continuous and discrete concept representations. Moreover, NCB features a straightforward inspection of learned concepts and facilitates easy revision procedures to align the obtained concept knowledge, e.g., with prior knowledge. In our evaluations, we show that NCB’s discrete concept-slot encodings retain the expressiveness of their continuous counterparts. Furthermore, they can easily be used in downstream applications via symbolic and transparent neural computations. In this context we introduce our novel CLEVR-Sudoku dataset that represents a challenging visual puzzle requiring both perception and reasoning capabilities (cf. Fig.  4).
@@ 2 Related Work
Unsupervised visual concept learning is concerned with obtaining concept-level representations from unlabeled images [19]. Where some works have tackled this only for specific domains, e.g., to extract “teachable” concepts for chess [53] or manipulation concepts via videos of task demonstrations [32] others depend on object-level concept guidance from initial image segmentations [21] or “natural supervision” [41]. In contrast, Vedantam et al. [68] and Wüst et al. [71] consider learning higher-level, relational concepts, i.e., assuming basic-level concepts have been provided. Related to their task-supervised manner of learning concepts, Wang et al. [69] obtain concepts via the training signal from their image classification task, thereby focusing on image-region-based concepts. Recently, several works have also investigated the potential of leveraging the knowledge stored in large pretrained models, e.g., via combinations of large-language models and CLIP embeddings [72, 43] or weakly-supervised queries to a vision-language model [4]. These works contain some form of supervision via text, class labels, or on the prompt level. In contrast, in this work we learn unsupervised concepts on both object and factor levels, which remain inherently inspectable and revisable.
@@
The motivation for inherently inspectable and revisable concept representations is to allow human stakeholders to investigate and potentially revise a model’s internal concepts. A majority of research focuses on post-hoc approaches based on distilling concept knowledge from pretrained models [73, 16, 74, 48, 13, 17]. In contrast, Lage and Doshi-Velez [29] investigate how to learn inspectable concept representations via human feedback, focusing hereby on tabular data and higher-level concepts. Additionally, Stammer et al. [60] learn inherently inspectable visual concepts via weak supervision and a prototype-based binding mechanism. No single work develops inherently inspectable and revisable concept representations in the context of unsupervised visual learning.
@@
The different properties of discrete vs. continuous encodings is a vibrant research topic that is very relevant for learning suitable concept representations. Continuous encodings allow for easier and more flexible optimization and information binding [36, 55, 56, 3]. Discrete representations are, however, considered to be vital for understanding AI models [25] and for mitigating shortcut learning [59], but also benefitial for solving complex visual reasoning tasks [20, 57]. However, learning discrete representations via neural modules is a challenging problem [37, 18, 15, 65]. Where some works have focused on categorical-distribution-based discretization [2, 22, 39], others have investigated the potential of retrieval-based discretization of continuous encodings via different forms of inherent “codebooks” [64, 62]. Only a few have investigated how to explicitly bind semantic visual information to specific discrete representations [60]. Where previous works mainly focus on one of the two representation types, we see great potential in the recent trend of explicitly integrating both types of representations [12].
@@
Figure 2: The Neural Concept Binder (NCB) combines (i) continuous, block-slot encodings via slot-attention based image processing with (ii) discrete, concept-slot encodings via retrieval-based inference. The retrieval corpus (distilled from the block-slot encodings) allows for easy concept inspection and revision by human stakeholders. Moreover, the resulting concept-slot encodings can be easily integrated into complex downstream tasks.
@@ 3 Neural Concept Binder (NCB): Extracting Hard from Soft Concepts
Our proposed Neural Concept Binder (NCB) framework tackles the challenge of learning inspectable and revisable object-factor level concepts from unlabeled images by combining two key elements: (i) continuous representations via (block-)slot-attention [36] based image processing with (ii) discrete representations via retrieval-based inference. Fig.  2 provides an overview of NCB’s inference, training, concept inspection, and revision processes. Let us now formally introduce these processes.
In this work, we refer to a concept as "the label of a set of things that have something in common" [1]. This definition can be applied on different scales of a visual scene: on an image level (e.g., an image of a park), an object level (e.g., a tree vs. a bird) or an object-factor level (e.g., the color of a bird). Overall, we consider a set of unlabeled images $X:=(x_{1},\cdots,x_{N})\in\mathbb{R}^{N\times D}$ with $x_{i}\in\mathbb{R}^{D}$, $N\in\mathbb{N}$ and $D\in\mathbb{N}$ (for simplicity, we drop the image index notation in the following). Briefly, given an image $x$, NCB derives a symbolic representation, $c$, which expresses the concepts of the objects in the image, i.e., object-factor level concepts. Herefore, NCB infers a block-slot encoding, $z$, of the image and performs a retrieval-based discretization step to finally infer concept-slot encodings, $c$. We begin by introducing the inference procedure of NCB. We hereby assume that NCB’s components have already been trained and will introduce details of the training procedure subsequently.
@@ 3.1 Inferring Concept-Slot Representations
Obtaining Continuous Block-Slot-Encodings. Consider an image $x\in X$. The first component of NCB, the soft binder, is represented via a block-slot encoder, $g_{\theta}:x\rightarrow z\in\mathbb{R}^{N_{S}\times N_{B}\times D_{B}}$, where $g$ is parameterized by $\theta$ (unless required, this notation is omitted in the following). The soft binder transforms an input image into a latent, continuous block-slot representation, where $N_{S}$ represents the number of slots, $N_{B}$ the number of blocks per slot, and $D_{B}$ the dimension of a block. It hereby utilizes two important types of binding mechanisms: spatial and factor binding. Spatial binding ensures spatial modularity across the entire scene and is achieved via slot attention [36]. As a result, each object in the image is represented in a specific slot, $z_{i}$, for slot $i$. Factor binding, on the other hand, as introduced by [56] ensures that different object factors (e.g., object attributes like color) are separately encoded in different blocks of a slot encoding, i.e., $z_{i}^{j}$ for block $j$. These two binding mechanisms jointly perform valuable object- and factor-based image processing. Overall, the resulting block-slot encodings represent continuous object-centric encodings of an input image (cf. Fig.  2), with objects encoded in slots and object factors encoded in blocks within the slots.
@@
Obtaining Discrete Concept-Slot-Encodings. The role of NCB’s second processing component, the hard binder, is to transform these continuous block-slot encodings into expressive, yet discrete concept-slot encodings. Specifically, the hard binder is represented via a retrieval encoder, $f$, (cf. Fig.  2). It processes the block-slot encodings, $z$, into a set of discrete concept-slot encodings, $c$. We refer to its underlying binding mechanism as concept binding. In detail, $f$ represents a function $f_{\mathcal{R}}:z\rightarrow c\in\mathbb{N}^{N_{S}\times N_{B}}$ that is parameterized by a retrieval corpus $\mathcal{R}$. This retrieval corpus consists of a tuple of sets $\mathcal{R}:=[\mathcal{R}^{1},\cdots,\mathcal{R}^{N_{B}}]$ where, in turn, each $\mathcal{R}^{j}:=\{(\texttt{enc}_{l}^{j},v_{l}):l\in\{1,...,|R^{j}|\}\}$ contains tuples of encodings, $\texttt{enc}_{l}^{j}\in\mathbb{R}^{D_{B}}$, and corresponding discrete values, $v_{l}\in\{1,\cdots,N_{C}\}$. Importantly, $\texttt{enc}_{l}^{j}$ consists of a representative block encoding of a specific concept. Such a concept is determined in NCB’s training phase and represents a distinct group of similar block representations (cf. Fig.  2, formal details below). The value $v_{l}$ thus represents the symbol identifier that corresponds to the specific concept that $\texttt{enc}_{l}^{j}$ is associated with. Hereby, each block contains up to $N_{C}\in\mathbb{N}$ different concepts. To infer the concept symbol of a sample’s block-slot encoding, NCB compares $z_{i}^{j}$ to the encodings in the corresponding block’s retrieval corpus, $\mathcal{R}^{j}$, and selects the most likely concept. Specifically, given a distance metric $d(\cdot,\cdot)$ and the block-slot encoding, $z_{i}^{j}$, the selection function $s_{\mathcal{R}}:z_{i}^{j}\rightarrow l\in\mathbb{N}$ finds the index $l$ of the closest encoding in the retrieval corpus: $s_{\mathcal{R}}(z_{i}^{j})=\operatorname*{argmin}_{l}d(\texttt{enc}_{l}^{j},z_{i}^{j})$ such that $(\texttt{enc}_{l}^{j},v_{l})\in\mathcal{R}^{j}$. This results in the concept representation for slot $i$ and block $j$ denoted as $c_{i}^{j}:=v_{s_{\mathcal{R}}(z_{i}^{j})}$. For the slot $i$, we thus obtain $c_{i}:=[c_{i}^{1},\cdots,c_{i}^{N_{B}}]$. The final concept-slot encoding is given by $c:=f_{\mathcal{R}}(z)=[c_{1},\cdots,c_{N_{S}}]$. We further refer to Sec. A.1 for details on an alternative, top-k selection function. Let us now move on to NCB’s training procedure.
@@ 3.2 Unsupervised Concept Learning via NCB
The training procedure of the Neural Concept Binder is separated into two subsequent steps where we provide an overview here and details in Sec. A.2. We formally describe these steps using the pseudo-code in Alg.  1. The first step consists of optimizing the encoder, $g$, to provide object-factorised block-slot encodings. It is optimized for unsupervised image reconstruction based on the decoder model, $g^{\prime}_{\theta^{\prime}}:z\rightarrow\tilde{x}\in\mathbb{R}^{D}$ (cf. Fig.  2) and utilising a mean squared error loss: $L=L_{\mathrm{MSE}}(x,g^{\prime}(g(x)))$. The goal of NCB’s second training step is to obtain the retrieval corpus, $\mathcal{R}$. This procedure is based on obtaining an optimal clustering of block encodings via an unsupervised clustering model, $h$, and distilling the resulting information from $h$ into explicit representations in the retrieval corpus. Thus, for each block $j$ a clustering model $h_{\phi^{j}}$ (cf. Fig.  2) is fit to identify a (potentially overparameterised) optimal set of clusters within a set of block encodings. This is based on an unsupervised criteria, e.g., a density-based score [44], and results in $N_{C}\in\mathbb{N}$ clusters. Next, for each cluster, $v\in\{1,\cdots,N_{C}\}$, representative block encodings, $\texttt{enc}^{j}$, are extracted from $h$. Such an encoding represents either an averaged prototype or instance-based exemplar encoding. The corresponding tuples $(\texttt{enc}^{j},v)$ are explicitly stored in the retrieval corpus $\mathcal{R}^{j}$, and the final retrieval corpus consists of the set of individual corpora for each block, $R=[R^{1},\cdots,R^{N_{B}}]$. Through this training procedure, NCB learns to unsupervisedly categorize the object-factor information from the latent encoding space of the soft binder and stores this information in a structured, symbolic, and accessible way in the hard binder’s retrieval corpus. We refer to the resulting clusters of each block as NCB’s concepts. In the following we denote concepts with a capital letter for the block and a natural number for the category id, e.g., $A3$. We note that in practice, it is further possible to finetune the block-slot encoder, $g$, through supervision from the hard binder (cf. Fig.  2), e.g., once initial categories have been identified, and can be achieved via a standard supervised approach. Ultimately, this allows for dynamically finetuning NCB’s concept representations. Let us now introduce how human stakeholders can inspect and revise NCB’s learned concepts.
@@
3.3 Inspecting NCB’s Concepts
One key advantage of NCB’s concept representations is their inherent readability and inspectability. Specifically, NCB inherently enables: (i) implicit, (ii) comparative, (iii) interventional and (iv) similarity-based inspection (cf. Fig.  3). Where the first three aim at investigating NCB’s explicit, symbolic concept space (stored in $\mathcal{R}$), the last one aims at investigating its latent, continuous concept space (stored in $\theta$).
(i) Implicit inspection queries the model to provide a set of examples for a specific concept. Essentially, this answers the question "What are examples of this concept?". NCB answers this question in two ways: by providing samples from the retrieval corpus corresponding to exemplars of the concept or by identifying additional data samples belonging to the concept at hand.
(ii) Comparative inspection, on the other hand, allows comparing two specifically different concepts, e.g., "Why does this object depict concept H5 and not concept H1?". NCB hereby provides examples for both concepts for the user to compare and potentially identify dissimilar properties. Ultimately, this form of inspection allows to answer questions of the form "Why not …?" and represents a valuable tool for in-depth and targeted concept inspection.
(iii) Interventional inspection allows to answer questions such as "What if this object would have concept H1?" To answer this question, NCB utilizes its decoder $g^{\prime}$. Specifically, by swapping the block $z_{i}^{j}$ of a data sample’s block-slot encoding with that of a representative sample, $(\texttt{enc}_{l}^{j},v_{l})\in\mathcal{R}^{j}$, NCB can provide an interventional image reconstruction, from which the effect of the swapped concept can be observed. Ultimately, this form of inspection allows to answer important questions of the form "What if …?".

### v2

@@ Abstract
The challenge in object-based visual reasoning lies in generating concept representations that are both descriptive and distinct. Achieving this in an unsupervised manner requires human users to understand the model’s learned concepts and, if necessary, revise incorrect ones. To address this challenge, we introduce the Neural Concept Binder (NCB), a novel framework for deriving both discrete and continuous concept representations, which we refer to as “concept-slot encodings”. NCB employs two types of binding: “soft binding”, which leverages the recent SysBinder mechanism to obtain object-factor encodings, and subsequent “hard binding”, achieved through hierarchical clustering and retrieval-based inference. This enables obtaining expressive, discrete representations from unlabeled images. Moreover, the structured nature of NCB’s concept representations allows for intuitive inspection and the straightforward integration of external knowledge, such as human input or insights from other AI models like GPT-4. Additionally, we demonstrate that incorporating the hard binding mechanism preserves model performance while enabling seamless integration into both neural and symbolic modules for complex reasoning tasks. We validate the effectiveness of NCB through evaluations on our newly introduced CLEVR-Sudoku dataset. Code and data at: project page.{NoHyper}These authors share equal contribution.{NoHyper}Correspondence to: Wolfgang Stammer ¡wolfgang.stammer@cs.tu-darmstadt.de¿.
@@ 1 Introduction
An essential aspect of visual reasoning is obtaining a proper conceptual understanding of the world by learning visual concepts and processing these into a suitable representation (cf. Fig.  1). The majority of current machine learning (ML) approaches that focus on visual concept-based processing utilize forms of supervised [34, 68, 32, 84], weakly-supervised [41, 69, 49, 63, 8, 82] or text-guided [29] learning of concepts. These approaches all require some form of additional (prior) knowledge about the relevant domain. An attractive alternative, though much more challenging, is to learn concepts in an unsupervised fashion. This comes with several challenges: (i) learning an expressive concept representation without concept supervision is intrinsically difficult [40], and (ii) there is no guarantee that learned concepts align with general domain knowledge [36, 85, 9] and (iii) can therefore be utilized for complex downstream tasks. Moreover, (iv) to trust that the learned concept representations are reliable for high stakes scenarios [15], it is necessary to make the model’s concept representations human-inspectable and -revisable [68, 69, 31] (cf. Fig.  1 (left)).
Figure 1: Unsupervised learning of concepts for visual reasoning. (left) Models that learn concepts from unlabeled data require inspectable and revisable concept representations. (right) Concepts obtained from the Neural Concept Binder (NCB) can be utilized both in (interpretable) neural and symbolic computations.
@@
This work proposes the Neural Concept Binder (NCB) framework to learn expressive, yet inspectable and revisable, concepts from unlabeled data. NCB combines continuous encodings, obtained via block-slot-based soft-binding, with discrete concept representations, derived through retrieval-based hard-binding. NCB’s soft binding leverages the object-factor disentanglement capabilities of the recent SysBinder mechanism [65]. Subsequently, NCB’s hard binding mechanism utilizes HDBSCAN [12, 13] to cluster the continuous block-slot encodings, distilling a structured corpus of discrete concepts from these clusters. This corpus enables the retrieval of discrete concept representations during inference by matching the continuous encoding with the closest entries in the corpus. Thus, to address the challenges of unsupervised concept learning, NCB integrates the strengths of both continuous and discrete concept representations. Moreover, NCB enables straightforward concept inspection and facilitates easy revision procedures, allowing alignment of the learned concepts with prior knowledge. In our evaluations, we demonstrate that NCB’s discrete concept-slot encodings retain the expressiveness of their continuous counterparts. Moreover, they can be seamlessly integrated into downstream applications via symbolic and interpretable neural computations (cf. Fig.  1 (right)). In this context, we introduce our novel CLEVR-Sudoku dataset, which presents a challenging visual puzzle that requires both perception and reasoning capabilities (cf. Fig.  4).
@@ 2 Related Work
Unsupervised visual concept learning focuses on obtaining concept-level representations from unlabeled images [25]. Some works have tackled this only for specific domains, such as extracting “teachable” concepts for chess [62] or learning manipulation concepts from videos of task demonstrations [39]. Others rely on object-level concept guidance through initial image segmentations [27] or “natural supervision” [49]. In contrast, Vedantam et al. [77] and Wüst et al. [81] focus on learning higher-level relational concepts, i.e., assuming that basic-level concepts have already been provided. Several approaches learn concepts from the training signal of an image classification task [78, 1, 14, 38], often focusing on image-region-based concepts [22]. More recently, several works have explored leveraging the knowledge stored in large pretrained models, such as combining large language models with CLIP embeddings [82, 52] or using weakly-supervised queries to a vision-language model [8]. These approaches still rely on some form of supervision, whether through text, class labels, or prompts. In contrast, this work focuses on learning unsupervised concepts at both the object and factor levels, ensuring that these concepts remain inherently inspectable and revisable.
@@
The motivation for inherently inspectable and revisable concept representations is to allow human stakeholders to investigate and potentially revise a model’s internal concepts. Most research in this area focuses on post-hoc approaches that distill concept knowledge from pretrained models [83, 21, 57, 18, 23]. In contrast, Lage and Doshi-Velez [35] explore learning inspectable concept representations through human feedback, focusing on tabular data and higher-level concepts. Similarly, Stammer et al. [69] develop inherently inspectable visual concepts using weak supervision and a prototype-based binding mechanism. However, no existing work addresses the development of inherently inspectable and revisable concept representations in the context of unsupervised visual learning.
@@
The properties of discrete vs. continuous encodings are a vibrant research topic that is highly relevant to learning suitable concept representations. Continuous encodings allow for easier and more flexible optimization and information binding [43, 64, 65, 7]. However, discrete representations are considered essential for understanding AI models [31], mitigating shortcut learning [68, 3], and solving complex visual reasoning tasks [26, 66]. Despite their advantages, learning discrete representations through neural modules remains a challenging problem [44, 24, 20, 74]. While some works have focused on categorical-distribution-based discretization [4, 28, 46], others have explored retrieval-based discretization of continuous encodings using various forms of inherent “codebooks” [73, 71]. Only a few studies have explicitly addressed how to bind semantic visual information to specific discrete representations [69]. Whereas previous works typically emphasize one of the two representation types, we see great potential in the recent trend of explicitly integrating both discrete and continuous representations [17, 32, 84, 51].
@@
Figure 2: The Neural Concept Binder (NCB) combines continuous, block-slot encodings via slot-attention based image processing with discrete, concept-slot encodings via retrieval-based inference. The structured retrieval corpus (distilled from the block-slot encodings) allows for easy concept inspection and revision by human stakeholders. Moreover, the resulting concept-slot encodings can be easily integrated into complex downstream tasks.
@@ 3 Neural Concept Binder (NCB): Extracting Hard from Soft Concepts
In this work, we refer to a concept as "the label of a set of things that have something in common" [2]. This definition can be applied on different scales of a visual scene: on an image level (e.g., an image of a park), an object level (e.g., a tree vs. a bird) or an object-factor level (e.g., the color of a bird). Our proposed Neural Concept Binder (NCB) framework tackles the challenge of learning inspectable and revisable object-factor level concepts from unlabeled images by combining two key elements: (i) continuous representations via SysBinder’s block-slot-attention [65, 43] with (ii) discrete representations via retrieval-based inference. Fig.  2 provides an overview of NCB’s inference, training, concept inspection, and revision processes. Let us formally introduce these processes.
Overall, we consider a set of unlabeled images $X:=(x_{1},\cdots,x_{N})\in\mathbb{R}^{N\times D}$ with $x_{i}\in\mathbb{R}^{D}$, $N\in\mathbb{N}$ and $D\in\mathbb{N}$ (for simplicity, we drop the image index notation in the following). Briefly, given an image, x, NCB infers latent block-slot encodings, z, and performs a retrieval-based discretization step on z to infer concept-slot encodings, c. These express the concepts of the objects in the image, i.e., object-factor level concepts. We begin by introducing the inference procedure of NCB. We hereby assume that NCB’s components have already been trained and will introduce details of the training procedure subsequently.
@@ 3.1 Inferring Concept-Slot Representations
Obtaining Continuous Block-Slot-Encodings. Consider an image $x\in X$. The first component of NCB, the soft binder, is based on the systematic binding mechanism [65] and is represented by a block-slot encoder (cf. Fig.  2 (i)), $g_{\theta}:x\rightarrow z\in\mathbb{R}^{N_{S}\times N_{B}\times D_{B}}$, where $g$ is parameterized by $\theta$ (for simplicity, this notation is omitted in the following). The soft binder transforms an input image into a latent, continuous block-slot representation, where $N_{S}$ represents the number of slots, $N_{B}$ the number of blocks per slot, and $D_{B}$ the dimension of each block. The soft binder employs two key types of binding mechanisms: spatial and factor binding. Spatial binding ensures spatial modularity across the entire scene and is achieved through slot attention [43], allowing each object in the image to be represented in a specific slot, $z_{i}$. Factor binding, introduced by Singh et al. [65], ensures that different object factors (e.g., attributes like color) are encoded in separate blocks of a slot, i.e., $z_{i}^{j}$. These two binding mechanisms work together to perform object- and factor-based image processing. We refer to Sec. A.1 for additional details on both systematic (factor) binding and slot attention. Overall, the resulting block-slot encodings represent continuous, object-centric representations of the input image, with objects encoded in slots and object factors encoded within the blocks of those slots.
@@
Obtaining Discrete Concept-Slot-Encodings. The role of NCB’s second processing component, the hard binder, is to transform the continuous block-slot encodings into expressive, yet discrete concept-slot encodings. Specifically, the hard binder is represented by a retrieval encoder, $f$ (cf. Fig.  2 (v)), which processes the block-slot encodings, $z$, into a set of discrete concept-slot encodings, $c$. In detail, $f$ defines a function $f_{\mathcal{R}}:z\rightarrow c\in\mathbb{N}^{N_{S}\times N_{B}}$, parameterized by a retrieval corpus $\mathcal{R}$ (cf. Fig.  2 (iv)). This retrieval corpus consists of a tuple of sets $\mathcal{R}:=[\mathcal{R}^{1},\dots,\mathcal{R}^{N_{B}}]$, where each set $\mathcal{R}^{j}:=\{(\texttt{enc}_{l}^{j},v_{l}):l\in\{1,...,|\mathcal{R}^{j}|\}\}$ contains tuples of block encodings, $\texttt{enc}_{l}^{j}\in\mathbb{R}^{D_{B}}$, and corresponding discrete values, $v_{l}\in\{1,\cdots,N_{C}\}$. Importantly, $\texttt{enc}_{l}^{j}$ is a representative block encoding of a specific concept cluster, determined during NCB’s training phase (cf. Fig.  2 (iii), detailed below). $v_{l}$ serves as the symbol identifier for the concept cluster associated with $\texttt{enc}_{l}^{j}$. Each block can contain up to $N_{C}\in\mathbb{N}$ different concepts. To infer the concept symbol for a sample’s block-slot encoding, NCB compares $z_{i}^{j}$ with the encodings in the corresponding block’s retrieval corpus, $\mathcal{R}^{j}$, and selects the most fitting concept. Specifically, given a distance metric $d(\cdot,\cdot)$ and the block-slot encoding $z_{i}^{j}$, the selection function $s_{\mathcal{R}}:z_{i}^{j}\rightarrow l\in\mathbb{N}$ (Fig.  2 (v)) finds the index $l$ of the closest encoding in the retrieval corpus: $s_{\mathcal{R}}(z_{i}^{j})=\operatorname*{argmin}_{l}d(\texttt{enc}_{l}^{j},z_{i}^{j})$ such that $(\texttt{enc}_{l}^{j},v_{l})\in\mathcal{R}^{j}$. This results in the concept representation for slot $i$ and block $j$, denoted as $c_{i}^{j}:=v_{s_{\mathcal{R}}(z_{i}^{j})}$. For slot $i$, the full concept representation is denoted as $c_{i}:=[c_{i}^{1},\dots,c_{i}^{N_{B}}]$ and the final concept-slot encoding as $c:=f_{\mathcal{R}}(z)=[c_{1},\dots,c_{N_{S}}]$. We refer to Sec. A.2 for details on an alternative top-k selection function. We further note that NCB’s flexibility, in principle, allows also to utilize the continuous encodings of its soft binder (Fig.  2 dashed arrow) in case a downstream task requires it. Let us now move on to NCB’s training procedure.
@@ 3.2 Unsupervised Concept Learning via NCB
The training procedure of the Neural Concept Binder is separated into two subsequent steps where we provide an overview here and details in Sec. A.3. We formally describe these steps using the pseudo-code in Alg.  1. The first step consists of optimizing the encoder, $g$, to provide object-factorised block-slot encodings. It is optimized for unsupervised image reconstruction based on the decoder model, $g^{\prime}_{\theta^{\prime}}:z\rightarrow\tilde{x}\in\mathbb{R}^{D}$ (Fig.  2 (ii)) and utilizing a mean squared error loss: $L=L_{\mathrm{MSE}}(x,g^{\prime}(g(x)))$. The goal of NCB’s second training step is to obtain the retrieval corpus, $\mathcal{R}$. This procedure is based on obtaining an optimal clustering of block encodings via an unsupervised clustering model, $h$, and distilling the resulting information from $h$ into explicit representations in the retrieval corpus. For each block $j$ a clustering model $h_{\phi^{j}}$ (Fig.  2 (iii)) is fit to identify a potentially overparameterised set of clusters within a set of block encodings (based on an unsupervised criterion, e.g., a density-based score [53]), resulting in $N_{C}\in\mathbb{N}$ clusters. Next, for each cluster, $v\in\{1,\cdots,N_{C}\}$, representative block encodings, $\texttt{enc}^{j}$, are extracted from $h$. Such an encoding represents either an averaged prototype or instance-based exemplar encoding. The corresponding tuples $(\texttt{enc}^{j},v)$ are explicitly stored in the retrieval corpus $\mathcal{R}^{j}$ (Fig.  2 (iv)) where we use the index $l$ to identify specific encodings in $\mathcal{R}^{j}$, leading to $\mathcal{R}^{j}:=\{(\texttt{enc}_{l}^{j},v_{l}):l\in\{1,...,|\mathcal{R}^{j}|\}\}$. Thus, $\texttt{enc}_{l}^{j}$ represents one block encoding of $\mathcal{R}^{j}$ that has been assigned to cluster $v_{l}$. Finally, $\mathcal{R}=[\mathcal{R}^{1},\cdots,\mathcal{R}^{N_{B}}]$ represents the final retrieval corpus, i.e., the set of corpora for each block. Through this training procedure, NCB learns to unsupervisedly categorize the object-factor information from the latent encoding space of the soft binder and stores this information in a structured, symbolic, and accessible way in the hard binder’s retrieval corpus. We refer to the resulting clusters of each block as NCB’s concepts and denote concepts with a capital letter for the block and a natural number for the category id, e.g., $A3$. We note that in practice, it is further possible to finetune the block-slot encoder, $g$, through supervision from the hard binder (cf. gray arrow in Fig.  2), e.g., once initial categories have been identified, and can be achieved via a standard supervised approach. Ultimately, this allows for dynamically finetuning NCB’s concept representations. Let us now introduce how human stakeholders can inspect and revise NCB’s learned concepts.
@@
3.3 Inspecting and Revising NCB’s Concepts